OpenStudy (anonymous):

In triangle ABC, a=3,b=5,c=7, find the measure of angle c?

OpenStudy (anonymous):

i have all three sides, so there is no need for the pythagorean theorem so where do i got from there?

OpenStudy (anonymous):

use the law of sines

OpenStudy (anonymous):

im not familiar

OpenStudy (anonymous):

sorry.., i mean use the law of cosines

OpenStudy (anonymous):

how do you know which one to use, and what are they?

OpenStudy (anonymous):

C^2 = A^2 + B^2 + 2AB cos c

OpenStudy (anonymous):

please try.., and i'll see

OpenStudy (anonymous):

120?

OpenStudy (anonymous):

what did u get it?

OpenStudy (anonymous):

120

OpenStudy (anonymous):

am i right?

OpenStudy (anonymous):

good joob! ur right!

OpenStudy (anonymous):

great, can you just explain why you chose law of cos over law of sin?

OpenStudy (anonymous):

sorry..,the right answer is not 120

OpenStudy (anonymous):

i dont understand...i did it right :(

OpenStudy (anonymous):

-1/2=cos C=120

OpenStudy (anonymous):

C^2 = A^2 + B^2 + 2AB cos c 7^2 = 3^2 + 5^2 + (2x 3x 5 ) cos c 49 = 9 + 25 + 30 cos c 49 = 34 + 30 cos c 49 - 34 = 30 cos c 15 = 30 cos c 1 = 2 cos c 1/2 = cos c am i wrong?

OpenStudy (anonymous):

ur formula is wrong, its -2AB cos c

OpenStudy (anonymous):

oh yes.., i'm wrong, i'm sorry..,

OpenStudy (anonymous):

ur relly good job.., the answer is 120

OpenStudy (anonymous):

lol ok well.. why do we use law of cos and not the law of sin??

OpenStudy (anonymous):

is there a law of sin? not sure

OpenStudy (anonymous):

oh yes, there is a law of sin, so do you know why?

OpenStudy (anonymous):

ok i'll explain.., If there is no angle is known, it will definitely use the law of cosines

OpenStudy (anonymous):

what if there is one known angle?

OpenStudy (anonymous):

if one known angle and all of side is known, it will definitely use the law of sines

OpenStudy (anonymous):

ok, are you positive, i have an exam tomorrow, im taking your word for it lol, i hope your right, thanks again :)

OpenStudy (anonymous):

ur welcome..., good luck for ur exam..,

OpenStudy (anonymous):

please tell me the result of ur exam, ok :)

OpenStudy (anonymous):

thanks! and i will!

OpenStudy (anonymous):

@Life there is reference for u please read it deeply http://www.mathwarehouse.com/trigonometry/law-of-sines-and-cosines.php

OpenStudy (anonymous):

wow, thank you so much, this is great!

OpenStudy (anonymous):

ur welcome @Life